### v1

@@ 6 Conclusion
With the great success of large language models and their emerging abilities in various downstream tasks, vision language models started to revolutionise various vision tasks such as image captioning and visual question answering. However, these emerging abilities are a product of a wide training phases of almost every data available on the web. It is evident that this pretraining data contains various forms of personal information that the model can inadvertently memorize and reveal during deployment. This is a serious violation to people privacy and must be addressed.
In this work, we examine to what extent can VLMs leak personal identities and whether the anonymization of the data used in the vision language alignment and finetuning phases is sufficient to prevent models from recognizing people identities. We show that in spite of the data anonymization, VLMs are still capable of recognizing people identities. We study the effect of context and simple image anonymization techniques like blurring and show that those have little to no effect on privacy leakage. This paper calls for new research to investigate this behaviour and design methods to mitigate the uncovered privacy risks.
@@ Appendix 0.B All results
Model $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$ BLIP-2flan-t5-xl [20] $0.95\pm 0.12$ $1.01\pm 0.88$ $6.94\pm 0.02$ $4.49\pm 6.2$ $7.21\pm 0.0$   $R_{0}$ 0.96 1.59 6.91 6.08 7.2   $R_{1}$ 1.26 0.8 7.16 6.46 7.09   $R_{2}$ 0.84 0.19 6.96 1.88 7.24   $R_{3}$ 0.56 2.61 6.79 5.74 7.26   $R_{4}$ 1.43 0.19 6.98 0.77 7.24   $R_{5}$ 0.64 0.68 6.85 6.03 7.21 BLIP-2opt [20] $1.43\pm 0.49$ $2.85\pm 3.41$ $5.24\pm 3.75$ $3.47\pm 3.0$ $4.6\pm 4.99$   $R_{0}$ 2.32 4.3 6.84 4.82 6.8   $R_{1}$ 0.5 1.49 4.18 4.36 2.81   $R_{2}$ 1.85 2.37 6.58 4.33 4.31   $R_{3}$ 1.33 5.88 5.73 4.81 6.32   $R_{4}$ 1.81 1.77 1.79 1.19 1.22   $R_{5}$ 0.76 1.27 6.29 1.32 6.16 LLaVA-1.57B [24] $0.12\pm 0.0$ $0.06\pm 0.0$ $4.14\pm 8.99$ $3.59\pm 9.69$ $9.56\pm 0.31$   $R_{0}$ 0.14 0.06 3.1 3.26 10.08   $R_{1}$ 0.12 0.03 9.22 5.99 10.08   $R_{2}$ 0.15 0.06 4.8 3.71 9.47   $R_{3}$ 0.1 0.16 0.64 0.13 9.62   $R_{4}$ 0.16 0.03 5.05 0.37 8.56   $R_{5}$ 0.08 0.03 2.02 8.09 9.54 LLaVA-1.6mixtral-7B [24] $0.35\pm 0.05$ $0.18\pm 0.0$ $7.12\pm 1.74$ $1.5\pm 2.23$ $6.83\pm 1.02$   $R_{0}$ 0.25 0.23 7.8 1.37 7.97   $R_{1}$ 0.2 0.14 8.5 4.32 5.37   $R_{2}$ 0.26 0.21 7.93 1.74 6.41   $R_{3}$ 0.27 0.18 5.03 0.29 7.29   $R_{4}$ 0.81 0.14 7.44 0.89 6.16   $R_{5}$ 0.31 0.19 6.02 0.36 7.76 PaliGemma3b-mix-224 [5] $1.12\pm 1.98$ $2.79\pm 2.57$ $7.81\pm 0.73$ $6.18\pm 2.08$ $7.66\pm 0.7$   $R_{0}$ 0.03 1.96 7.61 5.37 7.84   $R_{1}$ 0.09 3.69 6.66 5.31 6.19   $R_{2}$ 1.68 1.67 8.39 6.48 8.49   $R_{3}$ 1.16 0.91 6.98 4.45 8.18   $R_{4}$ 0.1 3.14 8.5 7.04 8.04   $R_{5}$ 3.63 5.33 8.7 8.46 7.21

### v2

@@ 6 Conclusion
In this work, we examine to what extent can VLMs leak personal identities and whether the anonymization of the data used in the vision language alignment and finetuning phases is sufficient to prevent models from recognizing people identities. We show that, despite the data anonymization, VLMs are still capable of recognizing people identities. We study the effect of context and simple image anonymization techniques like blurring and show that those have little to no effect on privacy leakage. This paper calls for new research to investigate this behaviour and design methods to mitigate the uncovered privacy risks.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of HPC resources. This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU and supported by the EU project AI4TRUST (No.101070190).
@@ Appendix 0.B All results
Model $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$ BLIP-2flan-t5-xl [20] $0.948\pm$0.117 $1.011\pm$0.880 $6.943\pm$0.017 $4.494\pm$6.196 $7.208\pm$0.004   $R_{0}$ 0.957 1.586 6.912 6.081 7.201   $R_{1}$ 1.259 0.802 7.164 6.459 7.092   $R_{2}$ 0.839 0.193 6.958 1.880 7.239   $R_{3}$ 0.558 2.614 6.786 5.737 7.264   $R_{4}$ 1.431 0.189 6.983 0.772 7.243   $R_{5}$ 0.642 0.684 6.853 6.035 7.210 BLIP-2opt [20] $1.428\pm$0.487 $2.847\pm$3.410 $5.235\pm$3.747 $3.471\pm$2.999 $4.603\pm$4.994   $R_{0}$ 2.321 4.301 6.840 4.822 6.803   $R_{1}$ 0.504 1.494 4.184 4.364 2.808   $R_{2}$ 1.846 2.367 6.580 4.327 4.314   $R_{3}$ 1.330 5.884 5.728 4.809 6.316   $R_{4}$ 1.809 1.771 1.788 1.188 1.217   $R_{5}$ 0.755 1.267 6.291 1.318 6.161 LLaVA-1.57B [24] $0.124\pm$0.001 $0.062\pm$0.003 $4.139\pm$8.991 $3.593\pm$9.692 $9.557\pm$0.311   $R_{0}$ 0.138 0.059 3.101 3.265 10.080   $R_{1}$ 0.122 0.025 9.224 5.993 10.076   $R_{2}$ 0.151 0.059 4.797 3.714 9.467   $R_{3}$ 0.105 0.164 0.638 0.126 9.623   $R_{4}$ 0.155 0.034 5.053 0.373 8.557   $R_{5}$ 0.076 0.029 2.019 8.087 9.539 LLaVA-1.6mixtral-7B [24] $0.352\pm$0.053 $0.182\pm$0.001 $7.120\pm$1.742 $1.496\pm$2.233 $6.827\pm$1.023   $R_{0}$ 0.248 0.231 7.801 1.372 7.965   $R_{1}$ 0.197 0.138 8.498 4.322 5.372   $R_{2}$ 0.264 0.214 7.932 1.737 6.412   $R_{3}$ 0.273 0.176 5.027 0.294 7.294   $R_{4}$ 0.814 0.143 7.441 0.890 6.161   $R_{5}$ 0.315 0.189 6.022 0.361 7.759 PaliGemma3b-mix-224 [5] $1.116\pm$1.978 $2.787\pm$2.573 $7.808\pm$0.729 $6.184\pm$2.079 $7.659\pm$0.700   $R_{0}$ 0.029 1.964 7.613 5.372 7.839   $R_{1}$ 0.092 3.693 6.664 5.313 6.194   $R_{2}$ 1.679 1.670 8.393 6.475 8.494   $R_{3}$ 1.158 0.915 6.979 4.453 8.183   $R_{4}$ 0.105 3.143 8.502 7.038 8.036   $R_{5}$ 3.630 5.334 8.695 8.456 7.206